solve using the elimations methods.
4x+y=5
2x-y=7

Answer by mananth(16946)

4x+y=5
2x-y=7
Add the two equations
4x+y+2x-y=5+7
6x=12
x=12/6
x=2
4x+y=5
4*2 +y=5
y=-3